OpenStudy (anonymous):

An acidophilus culture containing 150 bacteria doubles in population every hour. Write a function representing the bacteria population for every hour that it passes.

OpenStudy (anonymous):

doubles in one hour \[150\times 2\] doubles again in two hours \[150\times 2^2\]and again \[150\times 2^3\] in \(t\) hours you have \[150\times 2^t\]